# Tomographic redshift dipole: Testing the cosmological principle

### v2

@@ Abstract
11footnotetext: dasferreira.pedro@gmail.com22footnotetext: valerio.marra@me.com
Introduction
The Cosmic Microwave Background (CMB) provides a snapshot of the early universe, with its temperature anisotropies offering insights into cosmic structures and dynamics. The most prominent of these anisotropies is the dipole, approximately 100 times larger than the fluctuations seen across smaller scales. This dipole is traditionally attributed entirely to our proper motion relative to the CMB rest frame. Such interpretation leads to an inferred velocity $v=(369.82\pm 0.11)$ km/s in the direction $(l,b)=(264.021\pm 0.011,48.253\pm 0.005)^{\circ}$1, which is used in astronomy to convert observed redshifts into cosmological ones2.
Cosmology’s standard model, the $\Lambda$ cold dark matter ($\Lambda$CDM) model, is founded on the cosmological principle, which asserts that the universe is statistically homogeneous and isotropic on scales larger than approximately 100 Mpc3, 4. Consequently, the primary source of the dipole observed in distant objects should be attributed to our peculiar velocity $\boldsymbol{v}$ relative to the CMB rest frame. This hypothesis underlies the Ellis-Baldwin test5, which posits that velocity estimates derived from different sources, such as x-ray6, radio galaxies (RG)7, 8, 9, 10, 11, 12, 13, 14, 15, quasars (QSO)16, 14, supernovae Ia (SNe)17, 18, and even gravitational waves19, 20, should agree. However, over the past 25 years, several studies utilizing different RG and QSO catalogs have consistently found that, though the peculiar velocity is generally consistent in direction with the CMB dipole, its amplitude is in tension, with a value of $\lvert\boldsymbol{v}\rvert\sim 1000$ km/s, as reviewed in Fig. 1. This quarter-century-long puzzle poses a significant challenge to the $\Lambda$CDM model, with the tension regarding the CMB dipole now reaching a $5\sigma$ level14, 15, suggesting that the CMB rest frame is not shared by all matter in the observable universe. It warrants serious consideration, as any deviations from large-scale homogeneity and isotropy, if confirmed, would fundamentally alter our understanding of the cosmos and the primordial physics that set the universe’s initial conditions.
@@ Introduction
This work005005005005001,00010001{,}0001 , 0001,50015001{,}5001 , 5002,00020002{,}0002 , 000[eBOSS]    [BOSS]    [eBOSS+BOSS]    $\lvert\boldsymbol{v}\rvert$ (km/s)
@@
Figure 1: Velocity estimates from dipoles in chronological order with $1\sigma$ error bars. Top: previous results from x-ray, radio galaxies (RG), quasars (QSO) and SNe Ia, with corresponding references. Bottom: results of this work, combining several types of objects in the range $2.20.2<z<$ 2.2. Dashed line: the velocity obtained by the CMB dipole, considering the peculiar velocity hypothesis. Gray region: the $3\sigma$ velocity upper bound by ref. 21.
@@
The possibility exists that part of the CMB dipole might originate from intrinsic cosmological phenomena, such as primordial fluctuations on the last-scattering surface22, the presence of a large local void23, “tilted universe” scenarios24, 25, and specific inflationary models26. This suggests that the CMB dipole might not solely result from our peculiar velocity relative to the CMB rest frame, but could also include an intrinsic cosmological component, which could account for discrepancies between velocities estimated from the CMB and those derived from other astronomical sources. However, by following the footprints of aberration and Doppler in harmonic space27, ref. 21 put the first constraint on the intrinsic CMB dipole and the findings are in agreement with the kinematic interpretation supported by the $\Lambda$CDM model. Before ref. 21, all the previous results from Fig. 1 could accommodate the observed CMB dipole if one allows a fine-tuned scenario with a non-negligible intrinsic dipole in the opposite direction of our peculiar velocity. Such a situation implies a reduced total dipole, which would be wrongly interpreted as a smaller velocity. This is not more a possibility, as can be seen in Fig. 1 where we show the $3\sigma$ upper limit, updated according to ref. 12, on our peculiar velocity as determined by ref. 21 using Planck’s temperature and polarization data and assuming a non-zero intrinsic dipole component.
@@
Recent measurements using SNe data17, 18 reported velocities below those expected by the CMB dipole, suggesting that systematic effects or an unexpectedly large clustering dipole could have an important role in this tension. Supporting this perspective, ref. 28 showed that the number count dipole of the new Quaia catalog of QSOs from Gaia could align with the CMB dipole, depending on the galactic cut applied. Additionally, ref. 29 reported a radio dipole amplitude exceeding CMB predictions, with the unexpected characteristic of amplitude increasing inversely with frequency. For x-ray, we will need to wait for the next results using the eROSITA satellite data30 to see if they will converge to a smaller $\lvert\boldsymbol{v}\rvert$. Moreover, ref. 9 draws attention to the fact that to convert directly the measured dipoles to the frame velocity, i.e., to considering it purely kinetic, is a naive approximation that does not consider biases due mainly to shot noise and masking.
@@
$z^{\prime}(\boldsymbol{\hat{n}}^{\prime})=z(\boldsymbol{\hat{n}})\delta(\boldsymbol{\beta},\boldsymbol{\hat{n}}^{\prime})\;,$ (1)
@@
$\delta(\boldsymbol{\beta},\boldsymbol{\hat{n}}^{\prime})=\frac{\sqrt{1-\beta^{2}}}{(1+\boldsymbol{\beta}\cdot\boldsymbol{\hat{n}}^{\prime})}\;.$ (2)
@@
$\hat{\boldsymbol{n}}^{\prime}=\frac{\hat{\boldsymbol{n}}\cdot\hat{\boldsymbol{\beta}}+\beta}{1+\hat{\boldsymbol{n}}\cdot\boldsymbol{\beta}}\,\hat{\boldsymbol{\beta}}+\frac{[\hat{\boldsymbol{n}}-(\hat{\boldsymbol{n}}\cdot\hat{\boldsymbol{\beta}})\hat{\boldsymbol{\beta}}]}{\gamma(1+\boldsymbol{\beta}\cdot\hat{\boldsymbol{n}})}\;,$ (3)
such that, for an intrinsically isotropic redshift distribution in the sky, the redshift dipole is31
$\boldsymbol{\Delta_{1}}=-\boldsymbol{\beta}+\mathcal{O}(\beta^{2})\;.$ (4)
@@
An observer at rest relative to the Hubble flow can still observe a dipole anisotropy arising from the universe’s large-scale structure. This effect, known as the clustering dipole, becomes increasingly significant at lower redshifts ($z\ll 1$), as fluctuations in density and velocity are more pronounced at smaller scales. This intrinsic contribution must be added to Eq. 4, leading to
$\boldsymbol{\Delta_{1}}=-\boldsymbol{\beta}+\boldsymbol{\Delta_{1,\rm{int}}}+\mathcal{O}(\beta^{2})\;.$ (5)
@@
The clustering dipole’s main contribution comes from peculiar motions, and can be estimated by computing the bulk flow velocity, which is the expected root mean square velocity of all sources in a survey. We found that the clustering dipole is negligible for the dataset we consider, having a magnitude smaller than 2% of the expected $\beta$31. Therefore, from now on, we will simply adopt $\boldsymbol{\Delta_{1}}=-\boldsymbol{\beta}$.
Data
@@
We used tracers from 5 different catalogs from eBOSS32 and BOSS33 SDSS programs: i) the QSO DR16 eBOSS catalog in the range ${0.8\!<\!z\!<\!2.2}$, encompassing a fraction of sky $f_{\rm sky}=0.14$ with $343708$ objects; ii) the Luminous Red Galaxies DR16 eBOSS catalog (LRG eBOSS) in the range ${0.6\!<\!z\!<\!0.9}$, covering a $f_{\rm sky}=0.14$ with $163249$ objects; iii) the massive galaxies DR16 eBOSS catalog (CMASS eBOSS) in the range $0.750.6\!<\!z\!<\!$ 0.75, embracing $f_{\rm sky}=0.25$ and $193298$ objects; iv) the massive galaxies DR12 BOSS catalog (CMASS BOSS) in the range $0.60.4\!<\!z\!<\!$ 0.6, with $f_{\rm sky}=0.26$ and $686370$ objects; v) the low-redshift galaxies from the LOWZ DR12 BOSS catalog in the range $0.40.2\!<\!z\!<\!$ 0.4, with $f_{\rm sky}=0.26$ and $433791$ objects. See Fig. 2 for the catalogs’ redshift histograms and Extended Data Fig. 1 for the footprints. All these catalogs are divided into North Galactic Cap (NGC) and South Galactic Cap (SGC), and provide the weights to correct for imaging systematics, fiber collisions and redshift failures.
@@
Figure 2: Redshift distribution of SDSS data used. The bin size adopted is $\Delta z=0.001$, the same used in the estimator. Dashed lines indicate data constraints. From left to right, the first dashed line marks the lower bound, $z>0.2$, due to the lack of mocks below this value. The second dashed line indicates the upper bound for CMASS eBOSS data, $z<0.75$, implemented to prevent estimator misbehavior with sparsely populated bins. Similarly, for LRG data, we adopt $z<0.9$, denoted by the third dashed line.
@@
To infer the errors and biases of the estimator for eBOSS data, we used the 1000 EZmock realistic samples34, 35, which include effects of clustering, survey geometry, redshift evolution, sample selection biases, and all systematic effects. These samples were modified to incorporate the Doppler effect by applying Eq. (1). For the BOSS catalogs, we applied the same approach using 1000 MultiDark-Patchy mocks36. Both mock pipelines have been employed by the SDSS collaboration for comprehensive analysis of baryon acoustic oscillations and redshift space distortions37, 38, 39.
@@
The Emission Line Galaxy (ELG) eBOSS catalog was not considered as it covers merely $\sim 3\%$ of the sky. Given our aim to assess the largest scale modulation, the extent of sky coverage is crucial; hence, incorporating the ELG catalog yields negligible enhancement to our analysis. LOWZ objects with $z<0.2$ were not considered due to the lack of MultiDark-Patchy mocks for this redshift interval. Furthermore, CMASS eBOSS objects with $z>0.75$ were excluded, leading to a 4.6% reduction in the original catalog, and LRG eBOSS objects with $z>0.9$ were also omitted, accounting for a 6.6% reduction. These minor exclusions were necessary to prevent estimator misbehavior, as the estimator’s minimization becomes unstable in sparsely populated redshift bins (we adopted thin bins of $\Delta z=0.001$, see Methods).
Measuring the dipolar modulation
@@
Achieving an unbiased estimation of the dipole and a reliable assessment of its uncertainty presents a significant challenge, especially when targeting a subtle signal of $\Delta_{1}\sim 10^{-3}$. To tackle this issue, we introduce a novel dipole estimator, implement a comprehensive bias-removal pipeline, and quantify uncertainty using state-of-the-art SDSS mock catalogs. The key elements of our measurement technique are outlined below, with details available in the Methods section. Our dipole estimator performs a tomographic least squares fit of the Doppler modulation around the de-Dopplered redshift monopole – the rest-frame monopole recalculated to remove the expected Doppler effect – for each redshift bin of $\Delta z=0.001$, and separately for each galactic cap (NGC and SGC).
Figure 3: Distribution of monopoles ($\Delta_{0}$, average redshift) for the North (NGC) and South (SGC) Galactic Caps mocks. $\bar{\Delta}^{\rm NGC}_{0}$ represents the average NGC monopole. It is clear that the hemispheres have different depths.
First, separate analysis of the two galactic caps is essential for minimizing bias due to the differences between the NGC and SGC. Despite both caps being designed with equivalent equipment, analytical techniques, and observational strategies, a detailed examination of mocks and observational data reveal distinct depth variations between the hemispheres. As a result, the redshift monopoles $\Delta_{0}$ (average redshift considering the systematic weights) differ, as demonstrated in Fig. 3. These variations are comparable to, or exceed, the magnitude of the expected dipole $\Delta_{1}$ ($\sim 10^{-3}$). Given that the dipole is evaluated as a modulation around the monopole, acknowledging these discrepancies is essential to prevent bias in the combined estimation from the NGC and SGC datasets. Indeed, measuring the dipole, with and without hemisphere division, shows a typical deviation around $600$km/s for individual catalogs. When analyzing the two galactic caps together, the Doppler modulation is adjusted around the combined monopole of both hemispheres, with differences of the same order or bigger than the $\Delta_{1}c$ value of $\sim$370 km/s, expected within the $\Lambda$CDM model. Accounting for this effect is crucial for an accurate measurement.
@@
The second key point regards the estimation of the monopole itself. In a full-sky survey the monopole and de-Dopplered monopole are expected to be the same, as the Doppler modulation increases the redshifts in one direction by the same factor that decreases in the opposite direction. However, for any sky-limited survey this antipodal symmetry is broken, and one has to use the de-Dopplered monopole, that is, the original rest-frame monopole, in order to prevent the introduction of bias. Accordingly, our estimator considers objects individually, with their respective systematic weights, comparing the observed and expected boosted redshift for its coordinates, given the de-Dopplered monopole of the very narrow redshift bin and the velocity being tested. The final result is obtained by minimizing across all bins and hemispheres simultaneously.

### v3

@@ Abstract
1 Introduction
The Cosmic Microwave Background (CMB) provides a snapshot of the early universe, with its temperature anisotropies offering insights into cosmic structures and dynamics. The most prominent of these anisotropies is the dipole, approximately 100 times larger than the fluctuations seen across smaller scales. This dipole is traditionally attributed entirely to our proper motion relative to the CMB rest frame. Such interpretation leads to an inferred velocity $v=(369.82\pm 0.11)$ km/s in the direction $(l,b)=(264.021\pm 0.011,48.253\pm 0.005)^{\circ}$ [1], which is used in astronomy to convert observed redshifts into cosmological ones [2].
Cosmology’s standard model, the $\Lambda$ cold dark matter ($\Lambda$CDM) model, is founded on the cosmological principle, which asserts that the universe is statistically homogeneous and isotropic on scales larger than approximately 100 Mpc [3, 4]. Consequently, the primary source of the dipole observed in distant objects should be attributed to our peculiar velocity $\boldsymbol{v}$ relative to the CMB rest frame. This hypothesis underlies the Ellis-Baldwin test [5], which posits that velocity estimates derived from different sources, such as x-ray [6], radio galaxies (RG) [7, 8, 9, 10, 11, 12, 13, 14, 15], quasars (QSO) [16, 14], supernovae Ia (SNe) [17, 18], and even gravitational waves [19, 20], should agree. However, over the past 25 years, several studies utilizing different RG and QSO catalogs have consistently found that, though the peculiar velocity is generally consistent in direction with the CMB dipole, its amplitude is in tension, with a value of $\lvert\boldsymbol{v}\rvert\sim 1000$ km/s, as reviewed in Fig. 1. This quarter-century-long puzzle poses a significant challenge to the $\Lambda$CDM model, with the tension regarding the CMB dipole now reaching a $5\sigma$ level [14, 15], suggesting that the CMB rest frame is not shared by all matter in the observable universe. It warrants serious consideration, as any deviations from large-scale homogeneity and isotropy, if confirmed, would fundamentally alter our understanding of the cosmos and the primordial physics that set the universe’s initial conditions.
@@ 1 Introduction
This work002002002002004004004004006006006006008008008008001,00010001{,}0001 , 0001,20012001{,}2001 , 2001,40014001{,}4001 , 4001,60016001{,}6001 , 6001,80018001{,}8001 , 8002,00020002{,}0002 , 0002,20022002{,}2002 , 200              [eBOSS]                    [BOSS]      $\lvert\boldsymbol{v}\rvert$ (km/s)
@@
Figure 1: Velocity estimates from dipoles in chronological order with $1\sigma$ error bars. Top: previous results from x-ray, radio galaxies (RG), quasars (QSO) and SNe Ia, with corresponding references. Bottom: results of this work, combining several types of objects in the range $2.20.4<z<$ 2.2. Dashed line: the velocity obtained by the CMB dipole, considering the peculiar velocity hypothesis. Gray region and 2021 CMB result: the velocity measured by ref. 21 using CMB non-diagonal correlations and its $3\sigma$ upper bound.
@@
The possibility exists that part of the CMB dipole might originate from intrinsic cosmological phenomena, such as primordial fluctuations on the last-scattering surface [22], the presence of a large local void [23], “tilted universe” scenarios [24, 25], and specific inflationary models [26]. This suggests that the CMB dipole might not solely result from our peculiar velocity relative to the CMB rest frame, but could also include an intrinsic cosmological component, which could account for discrepancies between velocities estimated from the CMB and those derived from other astronomical sources. However, by following the footprints of aberration and Doppler in harmonic space [27], ref. 21 put the first constraint on the intrinsic CMB dipole and the findings are in agreement with the kinematic interpretation supported by the $\Lambda$CDM model. Before ref. 21, all the previous results from Fig. 1 could accommodate the observed CMB dipole if one allows a fine-tuned scenario with a non-negligible intrinsic dipole in the opposite direction of our peculiar velocity. Such a situation implies a reduced total dipole, which would be wrongly interpreted as a smaller velocity. This is not more a possibility, as can be seen in Fig. 1 where we show the $3\sigma$ upper limit, updated according to ref. 12, on our peculiar velocity as determined by ref. 21 using Planck’s temperature and polarization data and assuming a non-zero intrinsic dipole component.
@@
Recent measurements using SNe data [17, 18] reported velocities below those expected by the CMB dipole, suggesting that systematic effects or an unexpectedly large clustering dipole could have an important role in this tension. Supporting this perspective, ref. 28 showed that the number count dipole of the new Quaia catalog of QSOs from Gaia could align with the CMB dipole, depending on the galactic cut applied. Additionally, ref. 29 reported a radio dipole amplitude exceeding CMB predictions, with the unexpected characteristic of amplitude increasing inversely with frequency. For x-ray, we will need to wait for the next results using the eROSITA satellite data [30] to see if they will converge to a smaller $\lvert\boldsymbol{v}\rvert$. Moreover, ref. 9 draws attention to the fact that to convert directly the measured dipoles to the frame velocity, i.e., to considering it purely kinetic, is a naive approximation that does not consider biases due mainly to shot noise and masking.
@@ 2 The redshift dipole
$1+z^{\prime}(\boldsymbol{\hat{n}}^{\prime})=[1+z(\boldsymbol{\hat{n}})]\delta(\boldsymbol{\beta},\boldsymbol{\hat{n}}^{\prime})\;,$ (2.1)
@@
$\delta(\boldsymbol{\beta},\boldsymbol{\hat{n}}^{\prime})=\frac{\sqrt{1-\beta^{2}}}{(1+\boldsymbol{\beta}\cdot\boldsymbol{\hat{n}}^{\prime})}\;.$ (2.2)
@@
$\hat{\boldsymbol{n}}^{\prime}=\frac{\hat{\boldsymbol{n}}\cdot\hat{\boldsymbol{\beta}}+\beta}{1+\hat{\boldsymbol{n}}\cdot\boldsymbol{\beta}}\,\hat{\boldsymbol{\beta}}+\frac{[\hat{\boldsymbol{n}}-(\hat{\boldsymbol{n}}\cdot\hat{\boldsymbol{\beta}})\hat{\boldsymbol{\beta}}]}{\gamma(1+\boldsymbol{\beta}\cdot\hat{\boldsymbol{n}})}\;,$ (2.3)
such that, for an intrinsically isotropic redshift distribution in the sky, by expanding Eq. (2.1) in multipoles, the redshift dipole is [31]
$\boldsymbol{\Delta_{1}}=-\boldsymbol{\beta}+\mathcal{O}(\beta^{2})\;.$ (2.4)
@@
An observer at rest relative to the Hubble flow can still observe a dipole anisotropy arising from the universe’s large-scale structure. This effect, known as the clustering dipole, becomes increasingly significant at lower redshifts ($z\ll 1$), as fluctuations in density and velocity are more pronounced at smaller scales. This intrinsic contribution must be added to Eq. (2.4), leading to
$\boldsymbol{\Delta_{1}}=-\boldsymbol{\beta}+\boldsymbol{\Delta_{1,\rm{int}}}+\mathcal{O}(\beta^{2})\;.$ (2.5)
@@
The clustering dipole’s main contribution comes from peculiar motions, and can be estimated by computing the bulk flow velocity, which is the expected root mean square velocity of all sources in a survey. We found that the clustering dipole is negligible for the dataset we consider, having a magnitude smaller than 2% of the expected $\beta$ [31]. Therefore, from now on, we will simply adopt $\boldsymbol{\Delta_{1}}=-\boldsymbol{\beta}$, i.e., we disregard terms of the order $\mathcal{O}(\beta^{2})$.
3 Data
@@
We used tracers from 4 different catalogs from eBOSS [32] and BOSS [33] SDSS programs: i) the QSO DR16 eBOSS catalog in the range ${0.8\!<\!z\!<\!2.2}$, encompassing a fraction of sky $f_{\rm sky}=0.14$ with $343708$ objects; ii) the Luminous Red Galaxies DR16 eBOSS catalog (LRG eBOSS) in the range ${0.6\!<\!z\!<\!0.9}$, covering a $f_{\rm sky}=0.14$ with $163249$ objects; iii) the massive galaxies DR16 eBOSS catalog (CMASS eBOSS) in the range $0.750.6\!<\!z\!<\!$ 0.75, embracing $f_{\rm sky}=0.25$ and $193298$ objects; iv) the galaxies DR12 BOSS catalog in the range $0.60.4\!<\!z\!<\!$ 0.6 (mostly CMASS galaxies), with $f_{\rm sky}=0.26$ and $686370$ objects. See Fig. 2 for the catalogs’ redshift histograms and Fig. 3 for the footprints. All these catalogs are divided into North Galactic Cap (NGC) and South Galactic Cap (SGC), and provide the weights to correct for imaging systematics, fiber collisions and redshift failures.
@@
Figure 2: Redshift distribution of SDSS data used. The bin size adopted is $\Delta z=0.001$, the same used in the estimator. Dashed lines indicate data constraints. From left to right, the first dashed line indicates the upper bound for CMASS eBOSS data, $z<0.75$, implemented to prevent estimator misbehavior with sparsely populated bins. Similarly, for LRG data, we adopt $z<0.9$, denoted by the second dashed line.
Figure 3: Tracer distribution footprints. Number count maps for each tracer, presented in Mollweide projection using the HEALPix format with $N_{\rm side}=64$.
@@
To infer the errors and biases of the estimator for eBOSS data, we used the 1000 EZmock realistic samples [34, 35], which include effects of clustering, survey geometry, redshift evolution, sample selection biases, and all systematic effects. These samples were modified to incorporate the Doppler effect by applying Eq. (2.1). For the BOSS catalogs, we applied the same approach using 1000 MultiDark-Patchy mocks [36]. Both mock pipelines have been employed by the SDSS collaboration for comprehensive analysis of baryon acoustic oscillations and redshift space distortions [37, 38, 39].
@@
The Emission Line Galaxy (ELG) eBOSS catalog was not considered as it covers merely $\sim$3of the sky. Given our aim to assess the largest scale modulation, the extent of sky coverage is crucial; hence, incorporating the ELG catalog yields negligible enhancement to our analysis. BOSS objects with $z<0.4$ (mostly LOWZ galaxies) were not considered because it was not possible to minimize the estimator. We will come back to this issue in Section 5. Furthermore, CMASS eBOSS objects with $z>0.75$ were excluded, leading to a 4.6% reduction in the original catalog, and LRG eBOSS objects with $z>0.9$ were also omitted, accounting for a 6.6% reduction. These minor exclusions were necessary to prevent estimator misbehavior, as the estimator’s minimization becomes unstable in sparsely populated redshift bins (we adopted thin bins of $\Delta z=0.001$).
4 Measuring the dipolar modulation
4.1 Estimator
@@
Achieving an unbiased estimation of the dipole and a reliable assessment of its uncertainty presents a significant challenge, especially when targeting a subtle signal of $\Delta_{1}\sim 10^{-3}$. To tackle this issue, we introduce a novel dipole estimator, implement a comprehensive bias-removal pipeline, and quantify uncertainty using state-of-the-art SDSS mock catalogs. We will now discuss the estimator and the three key elements of our measurement technique.
We employ a least squares estimator to measure the dipole $\boldsymbol{\Delta}_{1}$ for each bin, based on the peculiar velocity hypothesis $\boldsymbol{\Delta}_{1}=-\boldsymbol{\beta}$. This calculation involves comparing the observed redshift distribution to the expected one for a theoretically isotropic $z$-field, boosted as per Eq. (2.1):
$\chi^{2}_{\rm bin}(\boldsymbol{\Delta}_{1})=\frac{\sum_{i}^{N}w_{i}\left[1+z_{i}-\Delta^{dD}_{0,{\rm bin}}\delta(-\boldsymbol{\Delta}_{1},\boldsymbol{\hat{n}}^{\prime}_{i})\right]^{2}}{\sum_{i}^{N}w_{i}},$ (4.1)
$w_{i,\mathrm{eBOSS}}$ $=w_{i,\mathrm{sys}}\cdot w_{i,\mathrm{noz}}\cdot w_{i,\mathrm{fc}}\,,$ (4.2)
$w_{i,\mathrm{BOSS}}$ $=w_{i,\mathrm{sys}}\cdot(w_{i,\mathrm{noz}}+w_{i,\mathrm{fc}}-1)\,,$ (4.3)
accounting for redshift failures ($w_{i,\mathrm{noz}}$), fiber collisions ($w_{i,\mathrm{fc}}$), and imaging systematics ($w_{i,\mathrm{sys}}$). Spectroscopic redshift measurement errors are negligible ($\sim 10^{-4}$[40, 33, 41] and thus not considered.
The first key point regards the estimation of the monopole itself. In a full-sky survey, the monopole and de-Dopplered monopole are expected to be identical because Doppler modulation symmetrically increases the redshifts in one direction while decreasing them in the opposite direction. However, this antipodal symmetry is disrupted in any sky-limited survey. To avoid introducing bias, it is necessary to use the “de-Dopplered” monopole $\Delta^{dD}_{0,{\rm bin}}$ of the respective bin, which is the average $1+z$ with the Doppler modulation removed,
$\Delta^{dD}_{0,{\rm bin}}=\sum_{i}^{N}w_{i}(1+z_{i})\delta(-\boldsymbol{\Delta}_{1},\boldsymbol{\hat{n}}^{\prime}_{i})^{-1}\bigg{/}\sum_{i}^{N}w_{i}\,,$ (4.4)
@@
that is, the original rest-frame monopole according to the tested value of $\boldsymbol{\Delta}_{1}$. Summarizing, our estimator considers objects individually, with their respective systematic weights, comparing the observed and expected boosted redshift for its coordinates, given the de-Dopplered monopole of the very narrow redshift bin and the velocity being tested. See Fig. 4 for an illustration of the estimation procedure.
Figure 4: Estimator diagram. Diagram of the $1+z$ spherical shell bin, its monopole and data. The estimator minimizes the difference between the observed $1+z_{i}$ (black points) and the monopole $\Delta^{dD}_{0,{\rm bin}}$ (the average of $1+z_{i}$ in the rest frame, described by the red line) times a dipole modulation $\delta(-\boldsymbol{\Delta}_{1},\boldsymbol{\hat{n}}^{\prime}_{i})$. The direction vector $\boldsymbol{\hat{n}}^{\prime}_{i}$ originates from the observer at the center of the shell.
Figure 5: Distribution of monopoles ($\Delta_{0}$, average redshift) for the North (NGC) and South (SGC) Galactic Caps mocks. $\bar{\Delta}^{\rm NGC}_{0}$ represents the average NGC monopole, i.e., average $1+z$. It is clear that the hemispheres have different depths.
The second key point concerns employing a separate analysis for the two galactic caps. Despite both caps being designed with equivalent equipment, analytical techniques, and observational strategies, a detailed examination of mocks and observational data reveal distinct depth variations between the hemispheres. As a result, the redshift monopoles $\Delta_{0}$ (average $1+z$ considering the systematic weights) differ, as demonstrated in Fig. 5. These variations are comparable to, or exceed, the magnitude of the expected dipole $\Delta_{1}$ ($\sim 10^{-3}$). Given that the dipole is evaluated as a modulation around the monopole, acknowledging these discrepancies is essential to prevent bias in the combined estimation from the NGC and SGC datasets. Indeed, measuring the dipole, with and without hemisphere division, shows a typical deviation around $600$ km/s for individual catalogs. When analyzing the two galactic caps together, the Doppler modulation is adjusted around the combined monopole of both hemispheres, with differences of the same order or bigger than the $\Delta_{1}c$ value of $\sim$370 km/s, expected within the $\Lambda$CDM model. Accounting for this effect is crucial for an accurate measurement. That is why we analyze objects in the Northern and Southern Galactic Caps separately. Therefore, to find the complete sample’s best-fit dipole $\boldsymbol{\Delta}_{1}$, we minimize across all bins according to:
$\chi^{2}(\boldsymbol{\Delta}_{\rm 1})=\sum_{\rm bin}\chi^{2}_{\rm bin,NGC}(\boldsymbol{\Delta}_{1})+\sum_{\rm bin}\chi^{2}_{\rm bin,SGC}(\boldsymbol{\Delta}_{1})\,.$ (4.5)
The final result is obtained by minimizing across all bins and hemispheres simultaneously.
4.2 Bias estimation and removal
The third key point of our analysis pipeline addresses estimating and removing potential biases in the dipole measurement. Systematic survey errors, including inconsistencies in observational strategies, footprint inhomogeneities, and variations in sky and telescope conditions, can introduce significant biases. To assess and correct for these biases, we utilized mock catalogs derived from eBOSS’s EZmock sample and BOSS’s MultiDark-Patchy mocks. Notably, the realistic version of EZmock incorporates observational systematics, a feature absent in the MultiDark-Patchy mocks, potentially affecting the findings related to the BOSS catalog.
For the BOSS mocks, which do not account for systematic effects as EZmocks do, we defined the weights for bias estimation as follows:
$w^{\rm{mocks}}_{i,\mathrm{BOSS}}=w_{i,\mathrm{veto}}\cdot w_{i,\mathrm{fc}},$ (4.6)
where $w_{i,\mathrm{veto}}$ serves as a flag, set to 1 for objects within the “veto mask” and 0 for those outside. The veto mask is designed to exclude sky sections not meeting specific quality standards, thereby minimizing areas heavily influenced by systematics. These standards may include regions with insufficient observational coverage, areas affected by contamination, or other conditions that could compromise data quality.
Figure 6: Probability distributions of de-biased results without boost. The points mark the $1\sigma$ confidence intervals.
To construct the bias map for both the eBOSS and BOSS surveys, we first Doppler boosted 100 mocks, randomly selected from the original pool of 1000 for each dataset, across 12 distinct directions identified by the pixel centers in the HEALPix $N_{\rm side}=1$ scheme [^1] HEALPix [42, 43], or Hierarchical Equal Area isoLatitude Pixelization, is an efficient system widely used in astrophysics and cosmology for pixelating the celestial sphere, enhancing the analysis and visualization of astronomical datasets.. These mocks were simulated at a velocity of 370 km/s, consistent with the kinematic interpretation of the CMB dipole. Our tests confirm that the bias map remains reliable across various dipole magnitudes. Additionally, we validated the accuracy of our estimator using the 4000 original mocks (1000 per catalog) without Doppler effects, successfully achieving expected results with bias correction, as depicted in Fig. 6.
For the bias vector calculation, we estimated the dipole for each of the 100 simulations per direction (1200 simulations per tracer). We computed the average difference vectors by subtracting the fiducial dipole vector $\boldsymbol{\Delta}^{\rm{fid,i}}_{1}$ for direction $i$ from the simulations’ average dipole vector $\langle\boldsymbol{\Delta}^{\rm{sim,i}}_{1}\rangle$. These vectors, expressed through their Cartesian components, define the pixel values of the three bias maps for each tracer. It is noteworthy that the detected bias is relatively minor. We performed bi-linear interpolation of these bias maps using the get_interp_val function from the healpy Python module [43]. This interpolation allows us to de-bias the measured dipole, ensuring precise analysis and interpretation. Fig. 7 illustrates this process for both combined and individual cases, displaying the average biased and de-biased dipole vectors against the benchmark dipoles from the simulations. The corrected bias for each Cartesian component aligns with the benchmark values, confirmed by comparing the mean difference vectors against the standard errors derived from 100 mocks per direction.